# FTFT: Efficient and Robust Fine-Tuning by Transferring Training Dynamics

### v2

@@ 1 Introduction
The problem of robustness can be mitigated using dual-model approaches. With such approaches, first a reference model is trained to estimate the importance of each training instance, and then a main model is trained based on the outputs of the reference model (Nam et al., 2020; Utama et al., 2020; Sanh et al., 2021; Karimi Mahabadi et al., 2020; Zhang et al., 2022; Liu et al., 2021). Of these approaches, dataset cartography (Swayamdipta et al., 2020) is especially attractive in view of its simplicity and the consistent improvements in robustness that it yields. It consists of three steps. First, a Data Map (DM) is constructed from the training dynamics (i.e. instance prediction probabilities) of a fine-tuning run of the reference model on the full dataset. This DM divides the training data into three subsets: ambiguous, hard-to-learn, and easy instances. Finally, the main model is fine-tuned using only either the ambiguous or hard-to-learn subset. An important question has to do with the choice of reference model. Swayamdipta et al. (2020) use the same PLM for both the reference and main model. However, a major drawback is the high computational cost, because it requires fine-tuning the same model twice. In this paper, we jointly address robustness and efficiency issues without sacrificing the simplicity of dataset cartography, by exploiting the transferability of training dynamics. We make three contributions.
First, we study the following question: Are DMs transferable across different model sizes and pretraining methods? We focus on the novel setting where DMs are constructed based on computationally efficient reference models to fine-tune more capable — and often larger — main models. The motivation is two-fold: 1) efficient reference models can ease the computational burden of constructing DMs, and 2) less capable reference models could be better at identifying ambiguous or hard training instances, because they are less likely to memorize training data (Tirumala et al., 2022; Carlini et al., 2023). We observe that, in most cases, training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2). We further show that the condition for successful transfers is a reasonably strong reference model, in a sense that we also seek to make precise (§4.3).
Second, we observe that fine-tuning using training instances selected by DMs enjoys consistently higher training efficiency than conventional fine-tuning (§6).
Third, building on these findings, we propose Fine-Tuning by transFerring Training dynamics (FTFT, §6): an efficient fine-tuning approach that leads to improved OOD performance. Compared to dataset cartography, FTFT uses more efficient reference models and early stopping. Experiments on two tasks, NLI and HSD, show that FTFT achieves better performance on OOD input than a standard approach based on Empirical Risk Minimization (ERM), while lowering the training cost by up to .
@@ 2 Background
Many studies have proposed dual-model approaches to improve model robustness that do not require knowledge of identifiable subsets of the data samples, such as NLI pairs in which hypotheses contain a negation (Gururangan et al., 2018), or HSD samples targetting a specific group (Dixon et al., 2018; Park et al., 2018). Nam et al. (2020) first train a reference model using generalized cross-entropy loss, and then train a main model while assigning higher weights to instances that are hard for the reference model. Sanh et al. (2021) use a Product-of-Expert (PoE) approach, by first training a reference model with limited capacity to capture dataset biases, and then training the main model to avoid these biases using PoE loss. Liu et al. (2021) propose the Just-Train-Twice approach (JTT), which involves first training a weak reference model using heavy regularization and vanilla SGD, and then up-weighing the training instances that the reference model predicts incorrectly when training the main model. Dataset cartography (Swayamdipta et al., 2020) is based on a similar idea, but it uses training dynamics instead of correctness to categorize training instances. We discuss this method below.
@@
Swayamdipta et al. (2020) use the same PLM as both the reference and the main model. In contrast, Sar-Shalom & Schwartz (2023) show that a DM constructed by (Clark et al., 2020) can be used to improve the robustness of (He et al., 2023). However, instead of using only the ambiguous subset, they added copies of this subset to the original training set to train the main model. Moreover, they did not investigate either DM transfer across model sizes and pretraining methods, or how such transferability can be exploited to improve efficiency.
@@
Our work is also connected to studies that have selected or reweighed data using reference models to improve in-distribution (ID) performance. Chang et al. (2017) use variance and proximity to the classification threshold from a reference model to reweigh training instances; Toneva et al. (2019) calculate the frequency of forgetting events (i.e. from correct to incorrect prediction), and remove the least forgettable instances; Paul et al. (2021) and Baldock et al. (2021) instead use error vector norm and effective prediction depth to estimate the contribution of a training instance.
@@ 3 Experimental Setup
We perform our experiments on two tasks, Natural Language Inference (NLI) and Hate Speech Detection (HSD). Following Swayamdipta et al. (2020), we set .
@@
To study model robustness, we include a few challenging OOD test sets for each task, in addition to the training set and an ID validation set. For NLI, we use the MultiNLI dataset (Williams et al., 2018) as the train and ID validation set, because of its diverse composition, covering 10 genres. We use AdversarialNLI (Nie et al., 2020) as the OOD test set, which consists of three rounds of adversarial data collection. AdversarialNLI is known to be challenging and it targets weaknesses of models trained on MultiNLI.
For HSD, we use CAD (Vidgen et al., 2021a) as the training and ID validation set. CAD consists of Reddit posts covering diverse topics and writing styles, annotated based on a fine-grained taxonomy. Following Ramponi & Tonelli (2022), we frame the task as a binary classification task, by marking identity-related abuse as hateful and other categories as non-hateful. As OOD test sets, we use DynaHate (Vidgen et al., 2021b), which contains three rounds of adversarial data collection and perturbations, and aligns with CAD’s hate speech definition (i.e. the datasets have the same labeling scheme).
@@
We mainly use DeBERTaV3 (He et al., 2023) and ELECTRA (Clark et al., 2020) in our experiments, because they offer strong performance and various model sizes. To study the transferability across different pretraining methods, we also use TinyBERT (Turc et al., 2020), BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as reference models.222 Costs for fine-tuning different PLMs are in Appendix A.1. We report FLOPs rather than GPU hours because we noticed occasional low GPU utilization especially when fine-tuning smaller PLMs. Besides ERM, we include a baseline that uses a random DM (i.e., random of the train data). We perform a search of optimal training steps on ERM, and use the same hyper-parameters for both ERM and data cartography approaches. Full details of the models and training setups are in Appendix A.1.
@@ 4 Transferability of Training Dynamics
We define successful transfers as transfers that produce comparable or better OOD performance than ERM. Our results show that, with a few exceptions, training dynamics are indeed transferable. To understand the conditions for successful transfers, we analyze the failure cases (§4.3), and find that the DMs of reference models that lead to successful transfers typically identify a larger subset as easy, which serves as a rough indicator of their capability. This finding can serve as a guideline for choosing reference models without training the main model, which is computationally expensive.
@@ 4.1 Transferability Across Model Sizes
The results for DeBERTaV3 are shown in Table 1 under section across different model sizes (NLI) and Table 5 (HSD, Appendix B). Training dynamics are transferable across different model sizes: when using as the main model, changing the reference model to either or yields comparable or even better performance. This observation is consistent with our hypothesis that efficient models are more sensitive to ambiguous and difficult instances. As a result, they can serve as more effective reference models compared to larger, more capable models. We also observe that, consistent with Swayamdipta et al. (2020), ERM achieves better ID performance (in the baselines section of Table 1), while data cartography performs comparably or better on OOD data. Note that hyper-parameters are determined by ERM performance (see Appendix A.1), making the training setup more favorable overall to ERM.
To investigate transferability of DMs further, we analyze whether reference models of different sizes identify similar groups of ambiguous instances. Figure 1(a) shows the percentage of ambiguous instances shared by reference models of different sizes and random seeds. We observe that the percentages of shared ambiguous instances between different sizes are only slightly smaller than those between the same size but different random seeds, providing further evidence for the transferability of DMs.333 The expected overlap between two random DMs is 0.33.
@@ 4.2 Transferability Across Pretraining Methods
Method Main Model Ref. Model Cost MultiNLI AdversarialNLI (Test) - R1 R2 R3 Baselines: ERM, ERM with Early Stopping, and Random DM ERM - 14.47 ERM - 28.29 ERM* - 100.00 ERM(ES) - 66.67 DM Random 100.00 Training Dynamics Transferability: Across Different Model Sizes DM 200.00 DM 114.47 DM 128.29 Training Dynamics Transferability: Across Different Pretraining Methods DM 104.61 DM 136.18 DM 216.78 DM 213.49 FTFT: Efficient Reference Models + Aggressive Early Stopping FTFT 51.97 FTFT 74.12 FTFT 79.93
We now study the transferability of training dynamics across different pretraining methods. Successful transfers of this type enable the exploitation of efficient reference models, in cases where there is no efficient version of the main model that suits the downstream task.
The results for as the main model with different reference models are shown in Table 1 under section across different pretraining methods (NLI) and Table 5 (HSD, Appendix B). We observe that training dynamics are generally transferable across different pretraining methods: achieves comparable performance using DMs constructed by different reference models in most cases. However, there are a few exceptions: when using as the reference model, the performance is clearly worse on the NLI OOD datasets than when using ERM. We hypothesize that is not strong enough for constructing effective DMs, and analyze this hypothesis further in §4.3 below.
4.3 How Efficient Can We Be?
Method Main Model Ref. Model Compute MultiNLI (Val.) AdversarialNLI (Test) - R1 R2 R3 ERM TinyBERT - 1.45 ERM - 4.61 ERM - 36.18 ERM - 14.47 ERM - 28.29 ERM - 113.49 ERM - 116.78 DM TinyBERT 101.45 DM 104.61 DM 136.18 DM 114.47 DM 128.29 DM 213.49 DM 216.78 DM TinyBERT 111.64 DM 114.80 DM 146.38 DM 124.67 DM 138.49 DM 223.68 DM 226.97
@@ 4.3 How Efficient Can We Be?
To answer this question, we compare the performance of a wide range of methods, see Table 2 (NLI) and Table 6 (HSD, in Appendix B). We include models from three categories: First, a wide range of seven models trained using ERM: TinyBERT, the small and base versions of DeBERTaV3 and ELECTRA, , and . Second, we use these ERM models as reference models to construct DMs, and use these DMs to fine-tune . By using reference models with different sizes and pretraining methods to fine-tune the same main model, we can inspect the impact of reference model capability on transferability. Third, we also include the results of using as the main model. By comparing results from different main models, we can better understand whether successful transfers originate from the compatibility between reference and main models, or the capability of the reference model itself. We make three observations.
First, weak reference models with poor ID performance, i.e., TinyBERT and , lead to unsuccessful transfers. However, the reference models do not need to be as capable as the main model: slightly weaker reference models could be more useful, indicated by the strong OOD performance of using as the reference model. Moreover, in these unsuccessful transfers, the main model OOD performance correlates with the reference model ID performance. For example, on NLI, TinyBERT has a lower ID performance than , and both main models and have a lower OOD performance when using TinyBERT as the reference model compared to when using as the reference model.
Second, whether a transfer is successful or not depends mostly on the reference model, rather than the compatibility between the reference and the main models: The transfers from TinyBERT and to both main models are unsuccessful, and the transfers from other reference models to both main models are successful.
@@
We now validate our reasoning. Given a reference model, we first split the training instances into two subsets based on mean : hard-to-learn ( of training instances) and other (the remaining ). We use a lower to make the difference clearer. Then, for each subset, we calculate the median in each epoch. We use median values because they are robust statistics of the central tendency. Figure 1(b) shows our results on MultiNLI, using two effective ( and ) and two ineffective ( and TinyBERT) reference models. We only include four models to make the figure clearer — we observe similar trends for other models: With effective reference models, hard-to-learn instances are gradually learned during training, while other instances already have high values from the first epoch. In contrast, with ineffective reference models, hard-to-learn data instances are not learned at all, suggested by their close-to-zero over different epochs; while other instances are gradually fitted, indicated by their increasing values.
To further validate our reasoning, we compute the percentages of training instances identified as easy by different reference models (Table 4 in Appendix B): We report results using to obtain more comprehensive results. Less effective reference models indeed identify fewer data points as easy. For example, on NLI with , TinyBERT identifies less than 20.0% of the instances as easy, compared to 46.65% by . Furthermore, the overlap between hard-to-learn and ambiguous instances in successful transfers is usually very large. For example, with , all effective reference models identify more than 46% as easy training instances (the maximum is 50%, when ambiguous and hard-to-learn data align perfectly).
@@ 5 Training Speed Gains From Data Selection
Our results from §4 suggest that one can improve the efficiency of dataset cartography with more efficient reference models. However, if we train the main models for the same number of steps as ERM, the cost of the full pipeline is still higher than conventional fine-tuning, because of the extra reference model training cost.
Previous studies have shown that training with a careful selection of “informative” training instances can lead to higher training speed than ERM (i.e. achieving the same performance with fewer training steps), despite the computational cost of selecting these instances being notoriously high (Sorscher et al., 2022; Feldman & Zhang, 2020; Paul et al., 2021; Toneva et al., 2019). Motivated by this, we now study whether we can achieve similar learning speed gains by training on instances selected by DMs. If such gains exist, we can further improve the efficiency of data cartography by training with fewer steps.
We present the performance of on HSD, fine-tuned using ERM and training instances selected by DMs, across different training durations (i.e. max training steps) in Figure 2. We include results for other datasets in Appendix B, on which we observe similar trends. Specifically, we make three observations. First, when the number of training steps is reduced, models fine-tuned with dataset cartography clearly outperform those fine-tuned with ERM, even in cases where ERM achieves better performance with more training steps (i.e., on ID validation data). Second, dataset cartography often yields better results with fewer training steps than with more, suggesting a tendency of overfitting in the late stage of training. Third, such efficiency gains hold consistently across different training durations. All three observations indicate that training with data instances selected by DMs achieves consistent higher training speed than ERM.

### v3

@@ Abstract
FTFT: Efficient and Robust Fine-Tuning by Transferring Training Dynamics
Yupei Du, Albert Gatt, and Dong Nguyen Utrecht University Utrecht, the Netherlands {y.du, a.gatt, d.p.nguyen}@uu.nl
@@ 1 Introduction
The problem of robustness can be mitigated using dual-model approaches. With such approaches, first a reference model is trained to estimate the importance of each training instance, and then a main model is trained based on the outputs of the reference model (Nam et al., 2020; Utama et al., 2020; Sanh et al., 2021; Karimi Mahabadi et al., 2020; Zhang et al., 2022; Liu et al., 2021). Among these approaches, dataset cartography (Swayamdipta et al., 2020) is especially attractive in view of its simplicity and the consistent improvements in model robustness. It consists of three steps. First, a Data Map (DM) is constructed from the training dynamics (i.e. instance prediction probabilities) of a fine-tuning run of the reference model on the full dataset. This DM divides the training data into three subsets: ambiguous, hard-to-learn, and easy instances. Finally, the main model is fine-tuned using only either the ambiguous or hard-to-learn subset. An important question has to do with the choice of reference model. Swayamdipta et al. (2020) use the same PLM for both the reference and main model. However, a major drawback of dataset cartography is the high computational cost, because it requires fine-tuning the same model twice. In this paper, we jointly address robustness and efficiency issues without sacrificing the simplicity of dataset cartography, by exploiting the transferability of training dynamics. We make three contributions.
First, we study the following question: Are DMs transferable across different model sizes and pretraining methods? We focus on the novel setting where DMs are constructed based on computationally efficient reference models to fine-tune more capable — and often larger — main models. Our motivation is two-fold: 1) efficient reference models ease the computational burden of constructing DMs, and 2) less capable reference models might be better at identifying ambiguous or hard training instances, because they are less likely to memorize training data (Tirumala et al., 2022; Carlini et al., 2023). Our results show that, in most cases, training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2). We further show that the condition for successful transfers is a reasonably strong reference model, which we also make precise in §4.3.
Second, we observe that fine-tuning with selected instances achieves consistently higher training efficiency than conventional fine-tuning (§6).
Third, building on these findings, we propose Fine-Tuning by transFerring Training dynamics (FTFT, §6): an efficient fine-tuning approach that leads to improved OOD performance. Compared to dataset cartography, FTFT uses efficient reference models and early stopping. Experiments on two tasks, NLI and HSD, show that FTFT achieves better performance on OOD input than conventional Empirical Risk Minimization (ERM), while lowering the training cost by up to .
@@ 2 Background
Many studies have proposed dual-model approaches to improve model robustness that do not require knowledge of identifiable subsets of the data samples, such as NLI pairs in which hypotheses contain a negation (Gururangan et al., 2018), or HSD samples targetting a specific group (Dixon et al., 2018; Park et al., 2018). Nam et al. (2020) first train a reference model using generalized cross-entropy loss, and then train a main model while assigning higher weights to instances that were hard for the reference model. Sanh et al. (2021) use a Product-of-Expert (PoE) approach, by first training a reference model with limited capacity to capture dataset biases, and then training the main model to avoid these biases using PoE loss. Liu et al. (2021) propose the Just-Train-Twice approach (JTT), which involves first training a weak reference model using heavy regularization and vanilla SGD, and then up-weighing the training instances that the reference model predicts incorrectly when training the main model. Dataset cartography (Swayamdipta et al., 2020) is based on a similar idea, but it uses training dynamics instead of correctness to categorize training instances. We discuss this method below.
@@
Swayamdipta et al. (2020) use the same PLM as both the reference and the main model. In contrast, Sar-Shalom and Schwartz (2023) show that a DM constructed by (Clark et al., 2020) can be used to improve the robustness of (He et al., 2023). However, instead of using only the ambiguous subset, they added copies of this subset to the original training set to train the main model. Moreover, they did not investigate either DM transfer across model sizes and pretraining methods, or how such transferability can be exploited to improve efficiency.
@@
Our work is also connected to studies that have selected or reweighed data using reference models to improve in-distribution (ID) performance. Chang et al. (2017) use variance and proximity to the classification threshold from a reference model to reweigh training instances; Toneva et al. (2019) calculate the frequency of forgetting events (i.e. from a correct to incorrect prediction), and remove the least forgettable instances; Paul et al. (2021) and Baldock et al. (2021) instead use error vector norm and effective prediction depth to estimate the contribution of a training instance.
@@ 3 Experimental Setup
We perform our experiments on two tasks, Natural Language Inference (NLI) and Hate Speech Detection (HSD). Following Swayamdipta et al. (2020), we select 33% the most ambiguous datapoints.
@@
To study model robustness, we include challenging OOD test sets for each task, in addition to the training set and an ID validation set. For NLI, we use the MultiNLI dataset (Williams et al., 2018) as the train and ID validation set, because of its diverse composition, covering 10 genres. We use AdversarialNLI (Nie et al., 2020) as the OOD test set, which consists of three rounds of adversarial data collection. AdversarialNLI is known to be challenging and it targets weaknesses of models trained on MultiNLI.
For HSD, we use CAD (Vidgen et al., 2021a) as the training and ID validation set. CAD consists of Reddit posts covering diverse topics and writing styles, annotated using a fine-grained taxonomy. Following Ramponi and Tonelli (2022), we frame it as a binary classification task, marking identity-related abuse as hateful and other categories as non-hateful. As OOD test sets, we use DynaHate (Vidgen et al., 2021b), since it aligns with CAD’s definition of hate speech. DynaHate contains three rounds of adversarial data collection and perturbations.
@@
We mainly use DeBERTaV3 (He et al., 2023) and ELECTRA (Clark et al., 2020) in our experiments, due to their strong performance and availability in multiple model sizes. To study the transferability across different pretraining methods, we also use TinyBERT (Turc et al., 2020), BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as reference models.222 Costs for fine-tuning different PLMs are in Appendix A.1. We report FLOPs rather than GPU hours because we noticed occasional low GPU utilization especially when fine-tuning smaller PLMs. Besides ERM, we include a baseline that uses a random DM (i.e., random of the training data). We perform a search of optimal training steps on ERM, and use the same hyper-parameters for both ERM and data cartography approaches. Full details of the models and training setups are in Appendix A.1.
Method Main Model Ref. Model Cost MultiNLI AdversarialNLI (Test) - R1 R2 R3 Baselines: ERM, ERM with Early Stopping, and Random DM ERM - 14.47 ERM - 28.29 ERM - 100.00 ERM(ES) - 66.67 DM Random 100.00 Training Dynamics Transferability: Across Different Model Sizes DM 200.00 DM 114.47 DM 128.29 Training Dynamics Transferability: Across Different Pretraining Methods DM 104.61 DM 136.18 DM 216.78 DM 213.49 FTFT: Efficient Reference Models + Aggressive Early Stopping FTFT 51.97 FTFT 74.12 FTFT 79.93
@@ 4 Transferability of Training Dynamics
We define successful transfers as transfers that produce comparable or better OOD performance than ERM. Our results show that, with a few exceptions, training dynamics are indeed transferable. To understand the conditions for successful transfers, we analyze the failure cases (§4.3). We find that the DMs of reference models that lead to successful transfers typically identify a larger subset as easy, which serves as a rough indicator of their capability. This finding can serve as a guideline for choosing reference models without training the main model, which is computationally expensive.
@@ 4.1 Transferability Across Model Sizes
The results for DeBERTaV3 are shown in Table 1 under section across different model sizes (NLI) and Table 5 (HSD, Appendix B). Training dynamics are transferable across different model sizes: when using as the main model, changing the reference model to either or yields comparable or even better performance. This observation is consistent with our hypothesis that efficient models are more sensitive to ambiguous and difficult instances. As a result, they can serve as alternative efficient reference models. We also observe that, consistent with Swayamdipta et al. (2020), ERM achieves better ID performance (in the baselines section of Table 1), while data cartography performs comparably or better on OOD data. Note that hyper-parameters are tuned for ERM performance (see Appendix A.1), making the training setup more favorable overall to ERM.
To investigate transferability of DMs further, we analyze whether reference models of different sizes identify similar groups of ambiguous instances. Figure 1(a) shows the percentage of ambiguous instances shared by reference models of different sizes and random seeds. The percentages shared between different model sizes are only slightly smaller than those between the same size but different random seeds, providing further evidence for the transferability of DMs.333 The expected overlap between two random DMs is 0.33.
@@ 4.2 Transferability Across Pretraining Methods
We now study the transferability of training dynamics across different pretraining methods. Successful transfers of this type enable the use of efficient reference models, when there is no efficient version of the main model that suits the downstream task.
The results for as the main model with different reference models are shown in Table 1 under section across different pretraining methods (NLI) and Table 5 (HSD, Appendix B). Training dynamics are generally transferable across different pretraining methods: achieves comparable performance using DMs constructed by different reference models in most cases. However, there is one exception: when using as the reference model, the performance is clearly worse on the NLI OOD datasets than when using ERM. We hypothesize that is not strong enough for constructing effective DMs. We analyze this hypothesis further in §4.3 below.
4.3 When Do Transfers Fail?
Method Main Model Ref. Model Compute MultiNLI (Val.) AdversarialNLI (Test) - R1 R2 R3 ERM Performance: Capabilities of Various Models ERM TinyBERT - 1.45 ERM - 4.61 ERM - 36.18 ERM - 14.47 ERM - 28.29 ERM - 113.49 ERM - 116.78 When Do Transfers Fail: Using Reference Models with Different Capabilities DM TinyBERT 101.45 DM 104.61 DM 136.18 DM 114.47 DM 128.29 DM 213.49 DM 216.78 DM TinyBERT 111.64 DM 114.80 DM 146.38 DM 124.67 DM 138.49 DM 223.68 DM 226.97
@@ 4.3 When Do Transfers Fail?
To answer this question, we compare the performance of a wide range of methods, see Table 2 (NLI) and Table 6 (HSD, in Appendix B). We include models from three categories: First, a wide range of seven models trained using ERM (section ERM Performance): TinyBERT, the small and base versions of DeBERTaV3 and ELECTRA, , and . Second, we use these ERM models as reference models to construct DMs, and use these DMs to fine-tune (section When Do Transfers Fail, first half). By using reference models with different sizes and pretraining methods to fine-tune the same main model, we can inspect the impact of reference model capability on transferability. Third, we also include the results of using as the main model (section When Do Transfers Fail, second half). By comparing results from different main models, we can better understand whether successful transfers originate from the compatibility between reference and main models, or the capability of the reference model itself. We make three observations.
First, weak reference models with poor ID performance, i.e., TinyBERT and , lead to unsuccessful transfers. However, the reference models do not need to be as capable as the main model: slightly weaker reference models could be more useful, indicated by the strong OOD performance when using as the reference model. Moreover, in these unsuccessful transfers, the main model OOD performance correlates with the reference model ID performance. For example, on NLI, TinyBERT has a lower ID performance than , and both main models and have a lower OOD performance when using TinyBERT as the reference model compared to when using as the reference model.
Second, whether a transfer is successful or not depends mostly on the reference model, rather than the compatibility between the reference and the main models: The transfers from TinyBERT and to both main models are unsuccessful, and the transfers from other reference models to both main models are successful.444 We also include results using as the main model, and TinyBERT and as reference models on HSD in Table 6. Transfers from substantially outperform ERM, while those from TinyBERT underperform ERM, further suggesting that the reference model is decisive for successful transfers.
@@
We now validate our reasoning. Given a reference model, we first split the training instances into two subsets based on mean : hard-to-learn ( of training instances) and other (the remaining ). We use a lower to make the difference clearer. Then, for each subset, we calculate the median in each epoch. We use median values because they are robust statistics of the central tendency. Figure 2 shows our results on MultiNLI, using two effective ( and ) and two ineffective ( and TinyBERT) reference models. We only include four models to make the figure clearer — we observe similar trends for other models: With effective reference models, hard-to-learn instances are gradually learned during training, while other instances already have high values from the first epoch. In contrast, with ineffective reference models, hard-to-learn data instances are not learned at all, suggested by their close-to-zero over different epochs; while other instances are gradually fitted, indicated by their increasing values.
To further validate our reasoning, we compute the percentages of training instances identified as easy by different reference models (Table 4 in Appendix B): We report results with . Less effective reference models indeed identify fewer data points as easy. For example, on NLI with , TinyBERT identifies less than 20.0% of the instances as easy, compared to 46.65% by . Furthermore, the overlap between hard-to-learn and ambiguous instances in successful transfers is usually high. For example, with , all effective reference models identify more than 46% as easy training instances (the maximum is 50%, when the ambiguous and hard-to-learn subsets overlap perfectly).
@@ 5 Training Speed Gains From Data Selection
Our results from §4 suggest that the efficiency of dataset cartography can be improved with more efficient reference models. However, if we train the main models for the same number of steps as ERM, the computational cost of the full pipeline is still higher than conventional fine-tuning (i.e., ERM), because of the extra reference model training cost.
Previous studies have shown that training with a careful selection of “informative" training instances can lead to higher training speed than ERM (i.e. achieving the same performance with fewer training steps), despite the computational cost of selecting these instances being notoriously high (Sorscher et al., 2022; Feldman and Zhang, 2020; Paul et al., 2021; Toneva et al., 2019). Motivated by this, we now study whether we can achieve similar learning speed gains by training on instances selected by DMs. If such gains exist, we can further improve the efficiency of dataset cartography by training with fewer steps.
We show the performance of on HSD, fine-tuned using the full training set (ERM) and using only the instances selected by DMs, across different training durations (i.e. max training steps) in Figure 1. Results on other datasets are in Appendix B, where we observe similar trends. We make three observations. First, when the number of training steps is reduced, models fine-tuned with dataset cartography clearly outperform those fine-tuned with ERM, even on ID validation data where ERM achieves better performance with more training steps. Second, dataset cartography often yields better results with fewer training steps than with more, suggesting a tendency of overfitting at the late training stage. Third, the efficiency gains hold consistently across different training durations. Our observations indicate that training with data instances selected by DMs consistently achieves higher training speed than ERM.